If a^2+b^2=11 and a*b=3 , find the value of (a-b)^2
Found 2 solutions by Edwin McCravy, richwmiller:
Answer by Edwin McCravy(20054)

If a squared plus b squared =11 and a times b=3 , find the value of (a-b)squared
       a² + b² = 11
            ab = 3

Multiply second equation through by -2

          -2ab = -6

Add equals to equals:

 a²       + b² = 11
     -2ab      = -6
 ------------------
 a² - 2ab + b² =  5

Factor the left side

(a - b)(a - b) = 5

      (a - b)² = 5

Edwin

Answer by richwmiller(17219)

(a-b)^2= a^2-2ba+b^2=11-6=5